The difference between two positive numbers is 5.Their product is 104.Find the numbers. solve and check.

Answer by edjones(8007)

x-y=5
xy=104
.
x=y+5
y(y+5)=104
y^2+5y=104
y^2+5y-104=0
(y+13)(y-8)=0
.
There are 2 possible answers:
y=-13, x=-8
or y=8, x=13
.
Ed